Parth (parthkohli):

Why is: \[\Large {x ^ {-m}} = {1 \over x^{m}}\] Note that x is not equal to 0. See below for the explanation.

Parth (parthkohli):

I'm gonna explain it to you in illustrations. MEDALS APPRECIATED. \[\Large \text{Illustration 1: }{{x^{x + 1} \over x} = x ^{x}}\] We have to agree, right? We know that when we divide two exponents with the same base, the powers subtract. Let's take an example of this. \[\Large {2^{3} \over 2} = 2^{2}} \] \[\Large \text{Illustration 2: } {x^{1} = x} \] We all know this one, right? \[\Large \text{Illustration 3: }{x^{0} = 1} \] Why is this so? See illustration 1 again and try to figure it out. (Hint, you can simply substitute x with 0.) \[\Large \text{Illutration 4:}{x^{-m} = {1 \over x^{m}}}\] Now, let us again see the illustration 1 and learn how it goes. Let's take an example of \[10^{-1}\] We know that \[x^{0} = 1\]. Inferring from example 1 and substituting x with -1, we can say: \[\Large {1 \over 10^{1}} = 10^{-1}\] \[\Large \text{ANOTHER WAY TO EXPLAIN}\] Taking an example of 5, we say that as the exponent decreases, we divide by 5. 1)25(which is 5 squared) divided by 5 = 5(which is 5 to the first power) 2)5(5 to the first power) divided by 5 = 1(5 to the zeroth power) 3)1(5 to the zeroth power) divided by 5 = 1 over 5(1 over 5 to the first power)

Parth (parthkohli):

\[\Large {2 ^ {3} \over 2} = 2^{2} \]

OpenStudy (anonymous):

\[x ^{m} \times x ^{-m}=1 that means x ^{m+(-m)}= x ^{0}=1\]

Parth (parthkohli):

A good example. But, that won't explain everything.

OpenStudy (anonymous):

get some good sort of qns

Parth (parthkohli):

This is a tutorial.

OpenStudy (anonymous):

u got ur ans or not finally

Parth (parthkohli):

THIS is a TUTORIAL, not a question.

OpenStudy (anonymous):

can u elaborate tutorial meaning pls

Parth (parthkohli):

Tutorial is an explanation of how to do something.

OpenStudy (anonymous):

operate x^m on both the sides and apply algebra

Parth (parthkohli):

I know, I'm explaining it WHY

OpenStudy (turingtest):

you should mention in your tutorial that this does not work for \(x=0\)

OpenStudy (phi):

I tried to explain it this way.